Solve the proportion
(c+3)/(C-1) = (c+2)/(c-3)

Answer by tutorcecilia(2152)

To begin, cross-multiply:
(c+3)(c-3) = (c-1)(c+2)
Multiply the factors using FOIL:
c^2 - 3c + 3c - 9 = c^2 + 2c - c - 2
Combine like terms:
c^2 - 9 = c^2 + - c - 2
Sovle for c:
c^2 - c^2 - 9 = c^2 -c^2 + - c - 2
- 9 = - c - 2
- 9 + 2 = - c - 2 + 2
-7 = c
Solve for c:
Checking the answer by plugging -7 back into the original equation:
(-7+3)/(-7-1) = (-7+2)/(-7-3)
-4/-8 = -5/-10
1/2 = 1/2 So, the answer checks out.